# CrossFi: A Cross Domain Wi-Fi Sensing Framework Based on Siamese Network

### v2

@@ Abstract
In recent years, Wi-Fi sensing has garnered significant attention due to its numerous benefits, such as privacy protection, low cost, and penetration ability. Extensive research has been conducted in this field, focusing on areas such as gesture recognition, people identification, and fall detection. However, many data-driven methods encounter challenges related to domain shift, where the model fails to perform well in environments different from the training data. One major factor contributing to this issue is the limited availability of Wi-Fi sensing datasets, which makes models learn excessive irrelevant information and over-fit to the training set. Unfortunately, collecting large-scale Wi-Fi sensing datasets across diverse scenarios is a challenging task. To address this problem, we propose CrossFi, a siamese network-based approach that excels in both in-domain scenario and cross-domain scenario, including few-shot, zero-shot scenarios, and even works in few-shot new-class scenario where testing set contains new categories. The core component of CrossFi is a sample-similarity calculation network called CSi-Net, which improves the structure of the siamese network by using an attention mechanism to capture similarity information, instead of simply calculating the distance or cosine similarity. Based on it, we develop an extra Weight-Net that can generate a template for each class, so that our CrossFi can work in different scenarios. Experimental results demonstrate that our CrossFi achieves state-of-the-art performance across various scenarios. In gesture recognition task, our CrossFi achieves an accuracy of 98.17% in in-domain scenario, 91.72% in one-shot cross-domain scenario, 64.81% in zero-shot cross-domain scenario, and 84.75% in one-shot new-class scenario. To facilitate future research, we will release the code for our model upon publication.
@@ I Introduction
Recently, Integrated Sensing and Communications (ISAC) has emerged as a prominent and popular technology direction aimed at enhancing the efficiency and intelligence of communication systems. Wi-Fi, as one of the key technologies in the realm of Internet of Things (IoT) communications, has found widespread application in various settings such as homes, offices, and public spaces [1]. In addition to its role in facilitating communication, Wi-Fi also holds promise as a sensing tool, owing to its characteristics including privacy protection, affordability, and penetration capability. In passive Wi-Fi sensing [2], by leveraging the variations in signal strength and multipath propagation caused by different objects and actions, it is possible to extract valuable information like Channel Statement Information (CSI) and Received Signal Strength Indicator (RSSI) to sense the environment and detect specific actions.
Wi-Fi sensing has attracted significant research attention, particularly in areas such as fall detection [3, 4], gesture recognition [5], and people identification [6]. These advancements have demonstrated the immense potential of Wi-Fi sensing in domains such as elderly care, military applications, and medical fields. However, a major challenge faced by existing Wi-Fi sensing models lies in their limited robustness. Even a slight change in the environment can lead to a significant deterioration in model performance or even complete failure. Addressing this critical issue is crucial for the practical deployment and utilization of Wi-Fi sensing devices.
@@
The rest of this paper is structured as follows: Section II introduces previous few-shot learning and zero-shot learning methods. Section III provides the basic principles of CSI and Wi-Fi sensing. Section IV introduces our method structure and system workflow according to different scenarios in detail. Section V presents comparative experiments and ablation studies to demonstrate the superiority of our method. Finally, Section VI concludes the paper and points to potential directions for further research.
@@ IV-C2 Training Phase
$l_{com}^{i,j}=$ $\alpha[\text{label}(c_{i})==\text{label}(c_{j}^{\prime})](1-S_{i,j})^{2}$
$+[\text{label}(c_{i})\neq\text{label}(c_{j}^{\prime})]S_{i,j}^{2}\ ,$
@@
where $c_{i},c_{j}^{\prime}$ represent the $i^{th}$ sample in the query branch and the $j^{th}$ sample in the key branch, $S_{i,j}$ represents the similarity between $c_{i}$ and $c_{j}^{\prime}$ calculated by the CSi-Net, $\text{label}(c_{i})$ indicates the category to which $c_{i}$ belongs, and $\alpha$ is a weight factor for positive pairs to solve the long-tail problem, as the number of positive pairs is usually significantly fewer than negative pairs.
@@
$l_{tem}^{i,j}=$ $\alpha[\text{label}(c_{i})==j](1-S_{i,j})^{2}$
$+[\text{label}(c_{i})\neq j]S_{i,j}^{2}\ ,$
@@
$s$ $=f(x,x,u)\ ,$ (6)
$t$ $=g(s,\theta_{t})\ ,$
$\hat{y}$ $=f(x^{\prime},t,v)\ ,$
$L$ $=l(\hat{y},y)\ ,$
$u$ $=v=\theta_{c}\ ,$
@@
$\frac{\partial L}{\partial\theta_{c}}$ $=\frac{\partial L}{\partial\hat{y}}(\frac{\partial\hat{y}}{\partial v}\frac{\partial v}{\partial\theta_{c}}+\frac{\partial\hat{y}}{\partial t}\frac{\partial t}{\partial s}\frac{\partial s}{\partial u}\frac{\partial u}{\partial\theta_{c}})$
$=\frac{\partial L}{\partial\hat{y}}(\frac{\partial\hat{y}}{\partial v}+\frac{\partial\hat{y}}{\partial t}\frac{\partial t}{\partial s}\frac{\partial s}{\partial u})\ ,$
@@ IV-D2 Few-shot
Figure 5: Effect of MK-MMD
@@
Figure 6: Domain Shift Problem.
@@ V-A1 Dataset
For the experiments, we utilized the dynamic part of the WiGesture Dataset [6]. The dataset was collected using an ESP32-S3 as the RX (receiver) and a home Wi-Fi router as the transmitter. The ESP device is equipped with an antenna and operates at a frequency of 2.4GHz with a sampling rate of 100 samples/s. The data collection environment is depicted in Fig. 8, with the transmitter and receiver positioned 1.5 meters apart. The dataset consists of a total of 8 individuals who performed various gestures, including left-right, forward-backward, and up-down motions, clockwise circling, clapping, and waving. These gestures are illustrated in Fig. 9.
@@ V-A2 Training Setup
In our experiments, we implement the CSi-Net and Weight-Net based on ResNet-18 [36]. The total number of parameters in our models is approximately 2.2 million. We observe that our model occupies around 1.5 GB of GPU memory when using a batch size of 64. For optimization, we use the Adam optimizer with an initial learning rate of $5\times 10^{-5}$ and a decaying rate of 0.01. Besides, our experiments are conducted on an Nvidia RTX 3090Ti GPU using the PyTorch framework.
@@ V-C Few-shot Experiment
Cross Domain New Class
Method Gesture Recognition People Identification Gesture Recognition People Identification
Siamese[13] 70.40% 82.87% 66.41% 80.92%
AutoFi (MLP-based)[10] 24.62% 24.71% 43.82% 81.75%
AutoFi (CNN-based)[10] 27.05% 36.14% 74.13% 86.58%
Yang et al.[5] 67.21% 74.22% 58.74% 49.00%
Ding et al.[19] 39.14% 59.50%
CrossFi w/ MK-MMD 91.72% 93.01% 80.62% 73.66%
CrossFi w/o MK-MMD 84.47% 87.50% 84.75% 81.97%
@@ V-E2 Effect of Multi-Attention Module
According to the results presented in Table V, the proposed attention measurement method demonstrates superior performance in most scenarios. However, it underperforms in the cross-domain one-shot scenario. As a result, we have decided to utilize Gaussian distance in few-shot scenario. It is worth noting that further investigation is warranted to uncover the underlying reasons behind this discrepancy and explore potential avenues for improvement.
@@ V-E3 Effect of Template Generation Method
VI Conclusion

### v4

@@ Abstract
In recent years, Wi-Fi sensing has garnered significant attention due to its numerous benefits, such as privacy protection, low cost, and penetration ability. Extensive research has been conducted in this field, focusing on areas such as gesture recognition, people identification, and fall detection. However, many data-driven methods encounter challenges related to domain shift, where the model fails to perform well in environments different from the training data. One major factor contributing to this issue is the limited availability of Wi-Fi sensing datasets, which makes models learn excessive irrelevant information and over-fit to the training set. Unfortunately, collecting large-scale Wi-Fi sensing datasets across diverse scenarios is a challenging task. To address this problem, we propose CrossFi, a siamese network-based approach that excels in both in-domain scenario and cross-domain scenario, including few-shot, zero-shot scenarios, and even works in few-shot new-class scenario where testing set contains new categories. The core component of CrossFi is a sample-similarity calculation network called CSi-Net, which improves the structure of the siamese network by using an attention mechanism to capture similarity information, instead of simply calculating the distance or cosine similarity. Based on it, we develop an extra Weight-Net that can generate a template for each class, so that our CrossFi can work in different scenarios. Experimental results demonstrate that our CrossFi achieves state-of-the-art performance across various scenarios. In gesture recognition task, our CrossFi achieves an accuracy of 98.17% in in-domain scenario, 91.72% in one-shot cross-domain scenario, 64.81% in zero-shot cross-domain scenario, and 84.75% in one-shot new-class scenario. The code for our model is publicly available at https://github.com/RS2002/CrossFi.
@@ I Introduction
Recently, Integrated Sensing and Communications (ISAC) has emerged as a prominent and popular technology direction aimed at enhancing the efficiency and intelligence of communication systems. Wi-Fi, as one of the key technologies in the realm of Internet of Things (IoT) communications, has found widespread application in various settings such as homes, offices, and public spaces [1]. In addition to its role in facilitating communication, Wi-Fi also holds promise as a sensing tool, owing to its characteristics including privacy protection, affordability, and penetration capability. In passive Wi-Fi sensing [2, 3], by leveraging the variations in signal strength and multipath propagation caused by different objects and actions, it is possible to extract valuable information like Channel Statement Information (CSI) and Received Signal Strength Indicator (RSSI) to sense the environment and detect specific actions.
Wi-Fi sensing has attracted significant research attention, particularly in areas such as fall detection [4, 5], gesture recognition [6, 7], and people localization [8, 9]. These advancements have demonstrated the immense potential of Wi-Fi sensing in domains such as elderly care, military applications, and medical fields. However, a major challenge faced by existing Wi-Fi sensing models lies in their limited robustness. Even a slight change in the environment can lead to a significant deterioration in model performance or even complete failure. Addressing this critical issue is crucial for the practical deployment and utilization of Wi-Fi sensing devices.
@@
The rest of this paper is structured as follows: Section II introduces previous few-shot learning and zero-shot learning methods. Section III provides the basic principles of CSI and Wi-Fi sensing. Section IV introduces our method structure and system workflow according to different scenarios in detail. Section V presents comparative experiments and ablation studies to demonstrate the superiority of our method. Section VI introduces a real-time system based on our method, highlighting its potential for practical Internet of Things (IoT) applications. Section VII examines the advantages and limitations of the proposed method. Finally, Section VIII concludes the paper and points to potential directions for further research.
@@ IV-C2 Training Phase
$l_{com}^{i,j}=$ $\alpha\mathbf{1}\{\text{label}(c_{i})=\text{label}(c_{j}^{\prime})\}(1-S_{i,j})^{2}$
$+\mathbf{1}\{\text{label}(c_{i})\neq\text{label}(c_{j}^{\prime})\}S_{i,j}^{2}\ ,$
@@
where $c_{i},c_{j}^{\prime}$ represent the $i^{th}$ sample in the query branch and the $j^{th}$ sample in the key branch, $S_{i,j}$ represents the similarity between $c_{i}$ and $c_{j}^{\prime}$ calculated by the CSi-Net, $\text{label}(c_{i})$ indicates the category to which $c_{i}$ belongs, $\alpha$ is a weight factor for positive pairs to solve the long-tail problem, as the number of positive pairs is usually significantly fewer than negative pairs, and the function $\mathbf{1}$ is the indicator function that returns 1 when the condition is true and 0 otherwise.
@@
$l_{tem}^{i,j}=$ $\alpha\mathbb{1}\{\text{label}(c_{i})=j\}(1-S_{i,j})^{2}$
$+\mathbb{1}\{\text{label}(c_{i})\neq j\}S_{i,j}^{2}\ ,$
@@
$s$ $=\text{f}(x,x;\theta_{c})\ ,$ (6)
$t$ $=\text{g}(s;\theta_{t})\ ,$
$\hat{y}$ $=\text{f}(x^{\prime},t;\theta_{c})\ ,$
$L$ $=\text{l}(\hat{y},y)\ ,$
@@
$\frac{\partial L}{\partial\theta_{c}}$ $=\frac{\partial L}{\partial\hat{y}}(\frac{\partial\text{f}(x^{\prime},t;\theta_{c})}{\partial\theta_{c}}+\frac{\partial\text{f}(x^{\prime},t;\theta_{c})}{\partial t}\frac{\partial\text{g}(s;\theta_{t})}{\partial s}\frac{\partial\text{f}(x,x;\theta_{c})}{\partial\theta_{c}})\ .$
@@ IV-D2 Few-shot
Figure 5: Effect of MK-MMD: The blue and red points represent samples from the source domain and target domain, respectively. The green line indicates the classification boundary. MK-MMD helps align the data distributions of the source and target domains in feature space, thereby enabling the classification boundary to function effectively in the target domain.
@@
Figure 6: Domain Shift Problem: This figure illustrates a potential failure caused by domain shift. If all the data in the target domain falls within the saturated region of the activation function, it renders the subsequent layers of the network ineffective (i.e. regardless of the inputs, the outputs will be similar).
@@ V-A1 Dataset
For the experiments, we utilized the dynamic part of the WiGesture Dataset [7] [^1] https://paperswithcode.com/dataset/wigesture. The dataset was collected using an ESP32-S3 as the RX (receiver) and a home Wi-Fi router as the transmitter. The ESP device is equipped with an antenna and operates at a frequency of 2.4GHz with a sampling rate of 100 samples/s. The data collection environment is depicted in Fig. 8, with the transmitter and receiver positioned 1.5 meters apart. The dataset consists of a total of 8 individuals who performed various gestures, including left-right, forward-backward, and up-down motions, clockwise circling, clapping, and waving. These gestures are illustrated in Fig. 9.
To further illustrate the cross-domain challenge present in the dataset, we employ t-Distributed Stochastic Neighbor Embedding (t-SNE) [46] to visualize the relationships between samples across different domains. As shown in Fig. 10, we observe that samples corresponding to the same person performing the same action tend to cluster closely together in the feature space. However, when the same action is performed by different people, the samples are often situated far apart, indicating significant variability in how different people execute the same gesture. This observation underscores the complexity of cross-domain recognition tasks, as the model must account for these variations to generalize effectively across different subjects.
Figure 10: t-SNE [46] Results for the WiGesture [7] Dataset: Different colors represent different gestures, while different shapes indicate different people.
@@ V-A2 Training Setup
In our experiments, we implement the CSi-Net and Weight-Net based on ResNet-18 [39]. The total number of parameters in our models is approximately 2.2 million. We observe that our model occupies around 1.5 GB of GPU memory when using a batch size of 64 (in-domain scenario). For optimization, we use the Adam optimizer with an initial learning rate of $5\times 10^{-5}$ and a decaying rate of 0.01. Besides, our experiments are conducted on an Nvidia RTX 3090Ti GPU using the PyTorch framework.
@@ V-C Few-shot Experiment
Cross Domain New Class Method Gesture Recognition People Identification Gesture Recognition People Identification Siamese[16] 70.40% 82.87% 66.41% 80.92% AutoFi (MLP-based)[13] 24.62% 24.71% 43.82% 81.75% AutoFi (CNN-based)[13] 27.05% 36.14% 74.13% 86.58% Yang et al.[6] 67.21% 74.22% 58.74% 49.00% Ding et al.[22] 39.14% 59.50% CrossFi w/ MK-MMD 91.72% 93.01% 80.62% 73.66% CrossFi w/o MK-MMD 84.47% 87.50% 84.75% 81.97%
@@ V-E2 Effect of Multi-Attention Module
According to the results presented in Table IX, the proposed attention measurement method demonstrates superior performance in most scenarios. However, it underperforms in the cross-domain one-shot scenario. Consequently, we have opted to utilize Gaussian distance in the series of experiments involving the WiGesture dataset in the few-shot scenario. It is worth noting that further investigation is warranted to uncover the underlying reasons behind this discrepancy and explore potential avenues for improvement. In the following sections, we will provide a preliminary explanation through experiments.
@@ V-E3 Effect of Template Generation Method
V-F Expanded Experiment
Figure 13: Cross-Domain Results in Expanded Datasets: The solid line represents CrossFi with the attention-based similarity calculator, the dashed line represents CrossFi with the Gaussian-based similarity calculator, and the dotted line indicates the benchmark results.
To further demonstrate the effectiveness of our model, we evaluate its cross-domain performance on additional datasets. Given that CrossFi is not exclusively designed for Wi-Fi sensing, we also conduct experiments on two image classification tasks. The datasets used are as follows:
  • $\bullet$
    Widar3.0 [19]: Widar3.0 is a large-scale Wi-Fi activity recognition dataset that encompasses multiple domains, including various individuals and rooms. Our experiments utilize data collected on November 9, 2018, which features six different gestures. We designate User1 as the source domain and User2 as the target domain. The dataset provides processed CSI known as the Body-coordinate Velocity Profile (BVP), with a dimension of $t\times 20\times 20$, where $t$ represents the time length. Following the approach in [19], we pad all samples to $t_{max}\times 20\times 20$ with zeros for standardization, and subsequently reshape them to $1\times 20t_{max}\times 20$ for processing by ResNet.
  • $\bullet$
    Digit-Five [51]: Digit-Five is a multi-domain handwritten digit recognition dataset that includes digits from 0 to 9. For our experiments, we select USPS as the source domain and MNIST as the target domain.
  • $\bullet$
    Office-Caltech10 [52]: Office-Caltech10 is a multi-domain visual object recognition dataset designed for office environments. It comprises 10 categories, such as books, laptops, and headphones. We choose Amazon as the source domain and Caltech as the target domain.
The results of the experiments are illustrated in Fig. 13, where the solid line represents CrossFi with the attention-based similarity calculator, the dashed line represents CrossFi with the Gaussian-based similarity calculator, and the dotted line indicates the benchmark result (zero-shot ResNet18). The results indicate that the proposed CrossFi demonstrates excellent performance across various datasets, particularly in Widar3.0 and Digit-Five, where it significantly outperforms the benchmark even in a zero-shot scenario. Additionally, we observe that in these three datasets, the Gaussian-based similarity calculator is not always better than attention-based method, which contrasts with our findings in the WiGesture dataset. Notably, in Widar3.0, the performance of the Gaussian-based method is not even monotonically increasing as the number of shots increases. We believe this is influenced by the extent of domain gap between the source domain and the target domain, which will be discussed in detail in Section VII.
VI IoT Deployment
To evaluate the practical applicability of our Wi-Fi sensing approach, we developed a real-time Internet of Things (IoT) system [^2] A demo is open source at: https://github.com/RS2002/ESP32-Realtime-System.. This system is designed to capture and process Wi-Fi signals continuously, enabling real-time monitoring and analysis.
VI-A Hardware Configuration
Our system, as shown in Fig. 14, is built around a Personal Computer (PC) that serves as the core unit for both training and deploying the inference models. Equipped with an NVIDIA RTX 3090 GPU, this robust hardware infrastructure supports the real-time processing requirements of our Wi-Fi sensing system, facilitating seamless data analysis and model inference.
Figure 14: Hardware: Our system consists primarily of a commercial Wi-Fi TX and an ESP32 RX. Additionally, a camera is employed to generate labels for the collected data during the training phase, while a remote PC receives the CSI forwarded by the ESP32 and is responsible for downstream calculations.
VI-B System Pipeline
The overall pipeline of the system is illustrated in Fig. 15, encompassing the entire process from data acquisition to inference. The pipeline begins with data collection, where the commercial Wi-Fi router sends ping reply packets to the ESP32-S3, simultaneously capturing CSI data and image frames via the network camera, following [4]. These data streams are transmitted to the PC for further processing.
During the training phase, the visual model generates corresponding labels from the captured images. These labels supervise the training of the CSI model (i.e. the proposed CrossFi), enabling it to predict environmental changes or object activities based on CSI data. Once the CSI model is adequately trained, the camera can be deactivated, and the system can transition to the inference phase. In this phase, the CSI model independently processes incoming CSI data to produce real-time sensing results without relying on visual inputs.
Figure 15: System Pipeline: The pipeline consists of two main phases: offline training and online inference. During the training phase, the ESP32 collects CSI while a camera generates labels for the collected samples. The aggregated data is then sent to a remote PC for training. In the inference phase, the ESP32 forwards the received CSI to the remote PC in real-time, allowing the PC to perform calculations based on the trained neural network.
VI-C Applications
Extensive testing has demonstrated that our system operates at a sampling rate of approximately 100 packets per second, which is sufficient to meet the requirements of various sensing tasks. One promising application involves integrating our Wi-Fi sensing system with existing smart home devices, such as refrigerators and air conditioners. By utilizing these devices as CSI receivers, the system can detect gestures or activities within a room and translate them into specific commands. This device-free remote operation offers a novel method for controlling smart home environments without the need for additional sensors, thereby enhancing both flexibility and user experience.
VII Discussion
VII-A Complexity and Performance
In the previous section, we introduced a real-time Wi-Fi sensing system based on CrossFi and discussed its practical deployment in IoT environments, where model complexity and scalability are critical considerations, as they directly impact real-time performance. However, in Section V-F, we observed a slight performance gap between our model and current state-of-the-art (SOTA) methods in image recognition tasks. A significant factor contributing to this gap is that the feature extractor used in CrossFi is ResNet18 [39], while most SOTA methods leverage larger models. Therefore, it is essential to strike a balance between model complexity and performance. Since the feature extractor component of CrossFi is not limited to ResNet18, we will compare different feature extractors in this section.
As shown in Table VII, we first compare the performance of various ResNet architectures as feature extractors. We observed that replacing ResNet18 with ResNet34 yields a notable improvement in model performance. However, when we use larger ResNet architectures, we find that model performance does not improve further and may even experience a slight decrease on the WiGesture dataset. One possible explanation for this is that the dataset may not be large enough; when the feature extractor is too large, it becomes challenging to learn the optimal representation of the data.
Next, based on ResNet18, we further explore the impact of using quantization and pruning methods. We found that these techniques can efficiently reduce model size with only a minimal decrease in performance. (It is important to note that the model size and GPU usage between the pruned and original methods remain the same due to the implementation in PyTorch, which masks the pruned parameters rather than deleting them.) As a result, we believe these methods can facilitate practical deployment. Additionally, we observed that even with ResNet101, inference can be completed within one second, processing approximately 350 samples. Thus, directly utilizing our proposed CrossFi method is also practical.
TABLE VII: Comparison of Complexity and Performance in One-Shot Cross-Domain Scenario: This method illustrates the feature extractor of CrossFi (with MK-MMD and attention-based similarity calculator). The quantization and pruning versions are based on ResNet18 [39].
Backbone Model Model Parameter Model Size GPU Occupation WiGesture [7] Office-Caltech10 [52]
ResNet34 [39] 4.26M 163.39MB 1.26GB 89.18% 89.05%
ResNet50 [39] 4.72M 180.96MB 2.36GB 87.36% 89.91%
ResNet101 [39] 8.53M 326.46MB 3.71GB 87.33% 89.07%
ResNet18 [39] 2.24M 85.66MB 0.93GB 80.42% 87.78%
Integer Quantization 2.24M 21.62MB 0.63GB 80.36% 84.03%
Pruning (20%) 1.79M 85.66MB 0.93GB 80.22% 84.81%
VII-B Key Components of CrossFi
The success of proposed CrossFi can be primarily attributed to its network architecture and training methodology. Comparing to a conventional Siamese network [16], the main innovations of CSi-Net and Weight-Net lie in the attention-based similarity calculator and the adaptive template generator. In this section, we will delve deeper into these two components to further elucidate their functionalities.
VII-B1 Adaptive Template Generator
In CrossFi, we propose an adaptive template generation method that utilizes Weight-Net to evaluate the quality of different samples. The quality score obtained from Weight-Net serves as the mixture ratio for template generation, allowing CSi-Net to be expanded into few-shot scenarios. In Section V-E3, we have already demonstrated its superior performance compared to other template generation methods. In this section, we conduct a small experiment to illustrate the efficiency of Weight-Net. After training the model, we select a batch of samples and input them into Weight-Net. We focus on one specific sample within this batch and add Gaussian noise, denoted as $N(0,\sigma^{2})$, to it. Table VIII presents the evaluated sample quality scores along with the standard deviation of the added noise. As $\sigma$ increases, the quality score exhibits a downward trend, indicating that the weight of this sample gradually decreases in the template. In this manner, Weight-Net effectively prioritizes high-quality samples over those affected by high noise.
TABLE VIII: Relationship Between Sample Quality Score and Added Gaussian Noise Standard Deviation
Gaussian Noise Variance 0 2 4 6 8 10
Sample Quality Score 0.4690 0.4673 0.4422 0.4474 0.4456 0.4278
VII-B2 Attention-based Similarity Calculator
In the experiments conducted, we found that the Gaussian-based and attention-based similarity calculators each have their own merits under different tasks. These two methods operate on different principles: the attention-based method functions as a pure black box, relying on neural networks to capture the similarity of samples, while the Gaussian-based method is highly interpretable, relying on the distance between the embedding outputs of samples. Intuitively, the Gaussian-based method depends on the similarity between the source domain and the target domain: only when these domains are similar are samples from the same category more likely to remain close in the embedding space. To demonstrate this, we conducted a series of one-shot experiments in a single source-domain scenario. We used person ID 0 as the source domain and person IDs 4 to 7 as the target domains, utilizing the WiGesture [7] dataset. As shown in Table IX, we compare the performance of the attention-based method with that of the Gaussian-based method. Additionally, to illustrate the similarity between the source domain and the target domain, we used the performance of ResNet18 trained in the source domain as a benchmark. A higher accuracy for the benchmark in the target domain indicates a higher similarity between the source and target domains. We observed that the performance gap between the Gaussian-based method and the attention-based method reflects the same trend as the benchmark performance (i.e., domain similarity). According to our results, we conclude that the attention-based method outperforms the Gaussian-based method when there is a significant gap between the source domain and the target domain. In the next step, we plan to combine the Gaussian-based and attention-based methods to address the problem under varying circumstances, leveraging the strengths of both approaches.
TABLE IX: One-shot Performance in a Single Source-Domain Scenario: The “Performance Gap” represents the accuracy difference between the Gaussian-based similarity calculator method and the attention-based method. The “Benchmark” refers to the zero-shot ResNet18 [39].
Source Domain: ID 0 Target Domain ID 4 6 7 5 Multi-Attention 94.79% 71.57% 93.29% 65.40% Gaussian Distance 82.18% 66.94% 96.87% 77.34% Performance Gap -12.61% -4.63% 3.58% 11.94% Benchmark [39] 19.51% 25.13% 31.82% 39.92%
VII-C Limitations and Future Directions
Although our CrossFi framework demonstrates excellent performance across various scenarios and tasks, we acknowledge that it has certain limitations that warrant further exploration in the future. First, in the zero-shot learning scenario, we currently select template samples from the target domains based on their similarity to source domain templates. However, this approach does not guarantee that the chosen target domain templates will be classified correctly. Nevertheless, the method remains efficient: even if a target domain template is misclassified, it still indicates a high degree of similarity to that category, and experimental results have shown promising performance. In the future, we aim to integrate other zero-shot learning methods with ours. If the target domain templates are accurately classified, the performance of the zero-shot learning method should converge with that of one-shot learning. Second, like most domain adaptation methods, our CrossFi currently supports only single target domain scenarios. To broaden its applicability, we can further develop a domain generalization approach based on our framework. Third, as mentioned in the previous section, the attention-based similarity calculation method has room for improvement. We can explore combining its strengths with Gaussian-based methods to enhance model performance further. Last but not least, this paper has demonstrated the model’s efficiency in wireless sensing and image recognition tasks. In the future, we can further explore the application of the model in other fields.
VIII Conclusion